if you have the corridents (1, 6) (4,r) and m=3 how do you find r
Answer by checkley77(12844)

(1, 6) (4,r) and m=3
SLOPE=(Y2-Y1)/X2-X1)
3=(R-6)/(4-1)
3=(R-6)/3
R-6=3*3
R-6=9
R=9+6
R=15 ans.
PROOF:
3=(15-6)/4-1)
3=9/3
3=3